OpenStudy (anonymous):

What is the volume of the solid obtained by rotating the region bounded by the curves y = x^3 - 5x^2 +6x ; y = 2x around the x-axis ?

OpenStudy (saifoo.khan):

Did you tried solving?

OpenStudy (saifoo.khan):

y = x^3 - 5x^2 +6x ; y = 2x First check where they both cut each other. x^3 - 5x^2 + 4x = 0 Solve for x. You should get: x = 0 , x =1 , x=4.

OpenStudy (saifoo.khan):

Next, Integrate the squared equation with applying limits. Vol of line - vol of curve.

OpenStudy (anonymous):

54PI/35

OpenStudy (anonymous):

is my correct

OpenStudy (anonymous):

*answer

OpenStudy (saifoo.khan):

Im getting 206.8

OpenStudy (anonymous):

i mean is that the final answer

OpenStudy (saifoo.khan):

Yes. My answer is final too.

OpenStudy (saifoo.khan):

Vol under curve = 2048pi/105 Vol under line = 256pi/3